OpenStudy (anonymous):

Willlll MEDAALLL!!!! Find the equation of the circle whose center and radius are given. center (0, 8), radius = 8

OpenStudy (sdfgsdfgs):

Standard form for circle eqn: https://www.mathsisfun.com/algebra/circle-equations.html

OpenStudy (anonymous):

I know the equation I just don't understand with how it is set up

OpenStudy (anonymous):

@Zweifel please help me

OpenStudy (anonymous):

Hi braedon

OpenStudy (anonymous):

please help me it's one where it iset up where I have to drag the items and I feel I don't have enough objects to drag

OpenStudy (anonymous):

\[r ^{2}=(x-h)+(y-k)\]

OpenStudy (anonymous):

yep i know

OpenStudy (anonymous):

it's in quiz 2 of coordinates

OpenStudy (anonymous):

what you do is put in the coordinates for h and k in the equation that I just sent and then you would square whatever they give you for a radius. That's basically it BTW h is the x coordinate and k is y that is all you have to do

OpenStudy (anonymous):

what you do is put in the coordinates for h and k in the equation that I just sent and then you would square whatever they give you for a radius. That's basically it BTW h is the x coordinate and k is y that is all you have to do

OpenStudy (anonymous):

\[64=(x-0)+(y-8)\]?

OpenStudy (anonymous):

missing some squares

OpenStudy (anonymous):

I thoughts so, but I was just following Zweifel

OpenStudy (anonymous):

Oops I forgot, eh you'll do fine person names B_______ n

OpenStudy (anonymous):

So I should have \[64=(x-0)^2+(y-8)^2\]

OpenStudy (anonymous):

Dude you already said my name once in here dingus!

OpenStudy (anonymous):

but really\[64=x^2+(y-8)\]

OpenStudy (anonymous):

Yes you are correct and hey where is my medal and how do you fan

OpenStudy (anonymous):

click on my user name and you'll see I button that says become a fan

OpenStudy (anonymous):

Check

OpenStudy (anonymous):

You still aren't my fan